The perimeter of a fence is 84 meters. If the width is 2 meters less than the length, find the width.
Answer by checkley79(3341)

P=2L+2W
W=L-2
84=2L+2(L-2)
84=2L+2L-4
4L=84+4
4L=88
L=88/4
L=22 METERS IS THE LENGTH.
W=22-2=20 METERS IS THE WIDTH.
PROOF:
84=2*22+2*20
84=44+40
84=84